OpenStudy (anonymous):

Factor completely: 10x5 + 4x4 + 8x3 Prime 2(5x5 + 2x4 + 4x3) 2x3(5x2 + 2x + 4) 2x(5x4 + 2x3 + 4x2)

OpenStudy (ivettef365):

From 10 /4 /8 the GCF = 2 From x^5 / x^4 / x^3 = x^3 (you get the smallest one)

OpenStudy (ivettef365):

Therefore, you have 2x^3 outside

OpenStudy (ivettef365):

Then it would be 2x^3(5x^2+2x+4)